# Adversarial Pruning: A Survey and Benchmark of Pruning Methods for Adversarial Robustness

### v1

@@ 3.2.2 B-Methods: Pruning before Training
B-methods search a sparse subnetwork which is then trained in isolation to match the dense performances, following the widely known Lottery Ticket Hypothesis (LTH [2]). In this case, the pipeline is summarized in two steps: pruning and training. The pruning step involves a search for the winning ticket, which consists of preliminary training and pruning of the network, resulting in a subnetwork that, trained in isolation, matches the dense performances. The training step, instead, consists in isolating the pruned network with the original starting random initialization (or, in some cases, initializing with a new random initialization [19]) and then training the resulting subnetwork from scratch. We specify that the watershed in labelling methods conforming to such pipeline as B-methods opposed to A-methods is their rationale, which posits that there exist subnetworks that, when trained in isolation, can match the performances of their dense counterpart. Therefore, we consider the main training step of the network to be applied on the subnetwork and, consequently, the pruning being applied before training. We thus consider the first step of the pipeline as a simple architecture search. Extending from [2], LTH has been validated on robust models by incorporating, starting from [19], robust objectives in both pruning and training steps.
Pruning step. The pruning step in B-methods is represented by a sparse architecture search. The procedure starts from a dense, randomly initialized model. The model is then trained for a given number of iterations to be iteratively pruned (IT) or pruned at the end of training in one-shot (1S). B-methods use a robust objective such as AT [12] within the search, leading to robust winning tickets [19, 51, 52, 53, 55], even in the early training stages [54].
Training step. The training step consists in training from scratch the subnetwork found in the first pruning step, yet starting from the original initialization of the dense model (or a new random one). While the majority of the APs classified as B-methods use AT as training procedure [19, 51, 52, 54, 55], the case of RST [53] stands out: in RST, the pruning stage consists in a mask search on the randomly initialized model that leads to a robust ticket. This ticket is considered to be robust from scratch and does not need further training.
@@ 3.2.3 D-Methods: Pruning during Training
D-methods represent end-to-end pipelines by jointly learning a mask used to prune the model and the networks’ parameters. Therefore, such methods need a single training run and a given robust objective, representing an efficient alternative to other methods. Following also seminal work [60], D-methods use a dynamic approach by pruning and regrowing parameters (i.e., restoring previously pruned parameters to prune new ones while keeping the same sparsity rate) during robust training. In this way, the sparsity constraint $s_{r}$ can be guaranteed throughout the training run by compensating pruning with regrowing (i.e., if a new parameter has to be pruned, a new one has to be regrown). Such procedure implies that D-methods decouple their pipeline from a classical 1S/IT notation.
@@ 3.3 Specifics
Creating a binary mask $\bm{m}$, depends upon a set of specifics that are associated with pruning: (i) the structure, which defines the kind of parameters $\theta_{i}$ that are removed as in Eq. 2 (e.g., single weights or entire filters) [3, 61]; (ii) the locality, which defines whether the desired sparsity rate $s_{r}$ is accomplished locally in each layer or globally [29]; and (iii) the pruning criteria, which is responsible of identifying the parameters to be removed, and has been designed in different ways to circumvent the problem of Eq. 1. While structure and locality are independent of the goal aspiring to reach in the sparse network (e.g., robustness for APs), the pruning criterion can be designed in different ways and has thus been geared towards preserving adversarial robustness in AP methods.
@@ 3.3.1 Structure (S vs US)
When removing single weights, pruning is said to be unstructured ($US$), while when removing entire structures (e.g. convolutional filters), pruning is referred to as structured ($S$[61, 7, 9, 29]. Following the mask creation described in Eq. 2, the parameters $\theta_{i}$ will be represented by single weights when $US$, and filters, channels, or kernels when $S$. In the literature, $S$ pruning is often considered more relevant than $US$, since setting to 0 entire structures allows creating a lighter architecture; $US$ pruning instead requires dedicated hardware to be correctly exploited [62]. In fact, removing entire structures from a dense network equals to having a network with reduced size, which can thus allow to directly exploit the size reduction. On the other hand, removing single weights enables a network with a sparse pattern, which can only yield a limited speedup on regular hardware. However, $US$ pruning enables a higher degree of flexibility and allows the preservation of higher performances than structured pruning when compared to the same level of sparsity. In addition, $US$ it is easier to implement and represents a highly useful mathematical prototype and test bench while getting increasing support for practical applications [62]. Among the considered AP methods, 21/26 implement $US$ and 13/26 implement $S$.
@@ 3.3.2 Locality (L vs G)
When pruning is applied locally ($L$), the desired $s_{r}$ is reached equally in each layer, while when applied globally ($G$), the sparsity $s_{r}$ is reached globally, thus treating the parameters in each layer equally and pruning them independently from their belonging layer. Therefore, when pruning is $L$, the rule used to prune the parameters (i.e., the pruning criterion of Sect. 3.3.3) is applied with the same $s_{r}$ in each layer, while when pruning is $G$ it is extended to the whole network. Consequently, while in $L$ pruning each layer has the same sparsity equal to $s_{r}$, in $G$ pruning each layer can have different sparsities that, summed up, satisfy the desired $s_{r}$ globally. Interestingly, methods such as HARP [20] and FlyingBird [54], that allow for different sparsities among layers and are thus labelled as $G$, control and optimize the sparsity rate of each layer. Similarly, DNR [59] extends from RADMM and allows layers to adjust their sparsity dynamically. Among the AP methods, a total of 11/26 methods implement local pruning, while 15/26 adopt a global approach.
@@ 3.3.3 Pruning Criteria
The pruning criterion defines a rule based on which a parameter is pruned or not. In essence, this translates to defining a rule based on which the set $\mathcal{C}$ in Eq. 2 is created. The most classical criterion is represented by the Lowest Weight Magnitude (LWM) based pruning [3], which removes the weights with the lowest magnitude based on the assumption that they tend to have the least harmful impact on the loss. An AP method, however, aims to prune the parameters that cause the least drop in robustness. Consequently, many of the works in Table 1 extended the problem of Eq. 1 to the adversarial case. However, independently from the objective, the sparsity constraint still makes the problem non-convex and combinatorial: this led AP methods to adopt multiple solutions circumventing this limitation in different ways while attempting to preserve robustness.
@@
SOlver-based Lowest Weight Magnitude (SOLWM). The SOLWM criterion represents AP methods whose pruning criterion’s primary goal is to first solve the constrained optimization problem imposed by sparsity (hence creating a robust model while satisfying the desired $s_{r}$) and then prune the parameters with the lowest magnitude. Among the collected AP methods, we identified multiple AP methods using the alternating direction method of multipliers (ADMM) as a solver for the constrained optimization problem [17, 59, 21]. ADMM is an optimization algorithm that, via the definition of an augmented Lagrangian, decomposes the main non-convex combinatorial problem into two sub-problems solved iteratively by SGD. In both RADMM [17] and PwoA [21], ADMM is applied on a pretrained network before finetuning, while in the modified formulation of DNR [59], the optimization is leveraged during training. Alternatively to ADMM, we identified Deadwooding [45] relaxing the constraints using Lagrangian multipliers and proposing a Lagrangian dual method to solve the optimization problem. When pruning using a SOLWM criterion, the training process designed to circumvent the constraint equals regularizing the parameters with a higher penalty when their contribution to the robust loss is higher (thus shrinking more their values). Therefore, as a final step, such methods prune the parameters with the lowest magnitude (hence, with LWM), or as often worded differently, they threshold the lowest weights.
@@
REgularization-based Lowest Weight Magnitude (RELWM). This criterion represents AP methods directly applying a regularization to the network parameters, which implicitly reduces their magnitude, and then pruning the weights with the lowest magnitude itself. By having a generally low weight magnitude, and then pruning with LWM, the harm induced by pruning on adversarial robustness is believed to be reduced. We identified a total of two AP methods adopting such an approach: RSR [40], uses an L1 penalty term (i.e., lasso) to encourage sparsity in the network in the adversarial pretraining step, and then prunes the weights with the least magnitude; TwinRep [56], reparameterizes the original network with the element-wise product of two matrices (namely, $\bm{\theta}=\bm{W}_{1}\otimes\bm{W}_{2}$) and learns both matrices during training. The overall effect of this procedure is similar to adding a weight decay factor, ultimately reducing the magnitude of the weights and thus facilitating pruning them with a lower impact on the overall performances.
@@
Least Importance Score (LIS). Pruning according to the LIS entails defining, for each parameter, a score based on which to prune. We define a score as an induced measure defining how important the parameter is for the overall network robustness, thus not considering directly accessible measures such as the weight magnitude for which naive pruning criteria are already defined. As shown in Table 1, LIS is the most common criterion for AP methods. However, the way in which importance is defined varies profoundly among AP methods. The approaches in HYDRA [16], HARP [20], InTrain [18], RST [53] and Heracles [35], all define a continuous mask (i.e., a set of importance scores of equal dimension to the parameters matrix) in an empirical risk minimization problem. The mask is then optimized using a robust objective, and the weights corresponding to the lower importance score (i.e., the ones with the least impact on adversarial robustness) are finally pruned. Other APs, formulate the importance differently: MAD [38] computes an estimate of the adversarial saliency for each parameter and prunes the least salient ones, as done for filters, similarly, in [47]; FSRP [50], prunes the filters based on the high-frequency components of their feature maps based on the rationale that adversarial examples mainly exist within such frequencies; similarly, in BNAP [42], the scaling factor of batch normalization layers is found to be representative of these components, being thus multiplied to the weights to define importance. Every AP method using LIS, just like any other, creates a binary mask based on the importance obtained and the desired $s_{r}$, which is then multiplied to the parameters to prune.
@@
Other pruning criterion. We identified multiple AP methods resorting to naive criteria such as LWM [33, 54]. Using the LWM approach is also particularly common for B-methods (pruning before training), where the pipeline for the ticket search is typically prioritized more than the criterion [51, 52, 54, 55]. We additionally find RFP [43] employing the Highest Gradient Magnitude (HGM) for filter pruning. Finally, differently from all other APs, BCS-P [57] optimizes the network with a negative log-posterior loss combining a sparsity prior with a robust training objective. Then, the network parameters are sampled from the posterior distribution. For regrowing in D-methods instead, the usual approach is to regrow based on the inverse of the pruning criteria, such as for InTrain [18]. Similarly, TwinRep [56] reparameterizes directly with the product of the two matrices, while BCS-P [57] simply resamples the parameters. FlyingBird [54] instead uses the Highest Gradient Magnitude (HGM), while DNR [59], following [60], uses normalized momentum.
@@ 4 Benchmarking Adversarial Pruning
In Sect. 3, we classified the AP methods based on pipeline and specifics, and then analyzed the core disparities. Upon a categorization of the methods, we took on the challenge of comparing each AP’s robustness while acknowledging the proposed taxonomy. However, our investigation revealed the impracticability of directly deriving such a comparison due to the many differences in the experimental setups of these evaluations. In fact, newly published APs are compared on equal setups only to a few of the “top-notch" AP methods, and while such analysis can be sufficient to assess the AP novelty, it is not enough to compare the robustness of each AP method (i.e., our goal). Furthermore, we find multiple issues associated with the adversarial evaluations conducted in these works, which represent a well-known problem in the field of adversarial ML, ultimately leading to overestimating the robustness of the models [22, 23]. All these obstacles are implicit limitations in the absence of a benchmark. In this section, we thus present our benchmark implementation and re-evaluate the pruned models produced by AP methods with a comparable experimental setup and fair adversarial setting. In this regard, we first bring to light, in Sect. 4.1, a number of problems related to the current AP evaluations, which are often found to be below par with respect to the SoA. We then present the benchmark experimental settings in Sect. 4.2 and re-evaluate the available AP methods in Sect. 4.4, where we conclude by discussing the re-evaluation results with particular attention to our taxonomy.

### v2

@@ 3.2.2 B-Methods: Pruning before Training
B-methods search a sparse subnetwork which is then trained in isolation to match the dense performances, following the widely known Lottery Ticket Hypothesis (LTH [2]). In this case, the pipeline is summarized in two steps: pruning and training. The pruning step involves a search for the winning ticket, which consists of preliminary training and pruning of the network, resulting in a subnetwork that, trained in isolation, matches the dense performances. The training step, instead, consists in isolating the pruned network with the original starting random initialization (or, in some cases, initializing with a new random initialization [19]) and then training the resulting subnetwork from scratch. We specify that the watershed in labeling methods conforming to such pipeline as B-methods opposed to A-methods is their rationale, which posits that there exist subnetworks that, when trained in isolation, can match the performances of their dense counterpart. Therefore, we consider the main training step of the network to be applied on the subnetwork and, consequently, the pruning being applied before training. We thus consider the first step of the pipeline as a simple architecture search. Extending from [2], LTH has been validated on robust models by incorporating, starting from [19], robust objectives in both pruning and training steps.
Pruning step. The pruning step in B-methods is represented by a sparse architecture search. The procedure starts from a dense, randomly initialized model. The model is then trained for a few iterations to be iteratively pruned (IT) or pruned after the iterations in one-shot (1S). B-methods use a robust objective such as AT [12] within the search, leading to robust winning tickets [19, 50, 51, 52, 54], even in the early training stages [53].
Training step. The training step consists in training from scratch the subnetwork found in the first pruning step, yet starting from the original initialization of the dense model (or a new random one). While the majority of the APs classified as B-methods use AT as a training procedure, the case of RST [52] stands out: in RST, the pruning stage consists in a mask search on the randomly initialized model that leads to a robust ticket. This ticket is considered to be robust from scratch and does not need further training.
@@ 3.2.3 D-Methods: Pruning during Training
D-methods represent end-to-end pipelines by jointly learning a mask used to prune the model and the networks’ parameters. Therefore, such methods need a single training run and a robust objective, representing an efficient compromise. Following also seminal work [59], D-methods use a dynamic approach by pruning and regrowing parameters (i.e., restoring previously pruned parameters to prune new ones while keeping the same sparsity rate) during robust training. In this way, the sparsity constraint $s_{r}$ can be guaranteed throughout the training run by compensating pruning with regrowing (i.e., if a new parameter has to be pruned, a new one has to be regrown). Such procedure implies that D-methods decouple their pipeline from a classical 1S/IT notation.
@@ 3.3 Specifics
Creating a binary mask $\bm{m}$, depends upon a set of specifics that are associated with pruning: (i) the structure, which defines the kind of parameters that are removed (e.g., single weights or entire filters) [3, 60]; (ii) the locality, which defines whether the desired sparsity rate $s_{r}$ is accomplished locally in each layer or globally [27]; and (iii) the pruning criteria, which is responsible for identifying the parameters to be removed, and has been designed in different ways to circumvent the problem of Eq. 1. While structure and locality are independent of the goal aspiring to reach in the sparse network (e.g., robustness for APs), the pruning criterion can be designed in different ways and has thus been geared towards preserving adversarial robustness in AP methods.
@@ 3.3.1 Structure (S vs US)
When removing single weights, pruning is said to be unstructured ($US$), while when removing entire structures (e.g. filters), pruning is referred to as structured ($S$[60, 7, 9, 27]. In the literature, $S$ pruning is often considered more relevant than $US$, since removing entire structures allows creating a lighter architecture; $US$ pruning instead requires dedicated hardware to be correctly exploited [61]. In fact, removing entire structures from a dense network equals directly reducing size, which can thus allow to be directly exploited. On the other hand, removing single weights enables a network with a sparse pattern, which can only yield a limited speedup on regular hardware. However, $US$ pruning enables a higher degree of flexibility and allows the preservation of higher performances than structured pruning when compared to the same level of sparsity. In addition, $US$ it is easier to implement and represents a highly useful mathematical prototype and test bench while getting increasing support for practical applications [61]. Among the considered AP methods, 21/26 implement $US$ and 13/26 implement $S$.
@@ 3.3.2 Locality (L vs G)
When pruning is applied locally ($L$), the desired $s_{r}$ is reached equally in each layer, while when applied globally ($G$), the sparsity $s_{r}$ is reached globally, thus treating the parameters in each layer equally and pruning them independently from their belonging layer. Therefore, when pruning is $L$, the rule used to prune the parameters (i.e., the pruning criterion of Sect. 3.3.3) is applied with the same $s_{r}$ in each layer, while when pruning is $G$ it is extended to the whole network. Consequently, while in $L$ pruning each layer has the same sparsity equal to $s_{r}$, in $G$ pruning each layer can have different sparsities that, summed up, satisfy the desired $s_{r}$ globally. Interestingly, methods such as HARP [20] and FlyingBird [53], that allow for different sparsities among layers and are thus labeled as $G$, control and optimize the sparsity rate of each layer. Similarly, DNR [58] extends from RADMM and allows layers to adjust their sparsity dynamically. Among the AP methods, a total of 11/26 methods implement local pruning, while 15/26 adopt a global approach.
@@ 3.3.3 Pruning Criteria
The pruning criterion defines a rule based on which a parameter is pruned or not. In essence, this translates to defining the algorithm $\mathcal{C}$. The most classical criterion is represented by the Lowest Weight Magnitude (LWM) based pruning [3], which removes the weights with the lowest magnitude based on the assumption that they tend to have the least harmful impact on the loss. An AP method, however, aims to prune the parameters that cause the least drop in robustness. Consequently, many of the work in Table 1 extended the problem of Eq. 1 to the adversarial case. However, independently from the objective, the sparsity constraint still makes the problem non-convex and combinatorial: this led AP methods to adopt multiple solutions circumventing this limitation in different ways while attempting to preserve robustness.
@@
SOlver-based Lowest Weight Magnitude (SOLWM). The SOLWM criterion represents AP methods whose pruning criterion’s primary goal is to first solve the robust constrained optimization problem imposed by sparsity (hence creating a robust model while satisfying the desired $s_{r}$) and then prune the parameters with the lowest magnitude. Many of the collected AP methods resort to the alternating direction method of multipliers (ADMM) as a solver for the constrained optimization problem [17, 58, 21]. ADMM is an optimization algorithm that, via the definition of an augmented Lagrangian, decomposes the main non-convex combinatorial problem into two sub-problems solved iteratively by SGD. In both RADMM [17] and PwoA [21], ADMM is applied on a robust pretrained network before finetuning, while in the modified formulation of DNR [58], the optimization is leveraged during training. Alternatively to ADMM, we identified Deadwooding [44] relaxing the constraints using Lagrangian multipliers and proposing a Lagrangian dual method to solve the optimization problem. When pruning using a SOLWM criterion, the training process designed to circumvent the constraint equals regularizing the parameters with a higher penalty when their contribution to the robust loss is higher (thus shrinking their values). Therefore, as a final step, such methods prune the parameters with the lowest magnitude (hence, with LWM).
@@
REgularization-based Lowest Weight Magnitude (RELWM). This criterion represents AP methods explicitly regularizing the network parameters, which reduces their magnitude, and then pruning the weights with the lowest magnitude itself. By having a generally low weight magnitude and pruning with LWM, the harm induced by pruning on adversarial robustness is believed to be reduced. We identified a total of two AP methods adopting such an approach: RSR [39] uses an L1 penalty term (i.e., lasso) to encourage sparsity in the network in the adversarial pretraining step, and then prunes the weights with the least magnitude; TwinRep [55], reparameterizes the original network with the element-wise product of two matrices (i.e., $\bm{\theta}=\bm{W}_{1}\odot\bm{W}_{2}$) and learns both matrices during training. The overall effect of this procedure is similar to adding a weight decay factor, i.e., reducing the magnitude of the weights and facilitating pruning them with a lower impact on the overall performance.
@@
Least Importance Score (LIS). Pruning, according to the LIS, entails defining, for each parameter, a score based on which to prune. We define a score as an induced measure defining how important the parameter is for the overall network robustness, thus not considering directly accessible measures such as the weight magnitude for which naive pruning criteria are already defined. As shown in Table 1, LIS is the most common criterion for AP methods. However, the way in which importance is defined varies profoundly among AP methods. The approaches in HYDRA [16], HARP [20], InTrain [18], RST [52] and Heracles [34], all define a continuous mask (i.e., a set of importance scores of equal dimension to the parameters matrix) in an empirical risk minimization problem. The mask is then optimized using a robust objective, and the weights corresponding to the lower importance score (i.e., the ones with the least impact on adversarial robustness) are finally pruned. Other APs, formulate the importance differently: MAD [37] computes an estimate of the adversarial saliency for each parameter and prunes the least salient ones, as done for filters, similarly, in [46]; FSRP [49], prunes the filters based on the high-frequency components of their feature maps based on the rationale that adversarial examples mainly exist within such frequencies; similarly, in BNAP [41], the scaling factor of batch normalization layers is found to be representative of these components, being thus multiplied to the weights to define importance. Every AP method using LIS creates a binary mask based on the importance obtained and the desired $s_{r}$, which is then multiplied to the parameters to prune.
@@
Other pruning criteria. We identified multiple AP methods resorting to naive criteria such as LWM [32, 53]. Using the LWM approach is also particularly common for B-methods (pruning before training), where the pipeline for the ticket search is typically prioritized more than the criterion, such as in [50, 51, 53, 54]. We additionally find RFP [42] employing the Highest Gradient Magnitude (HGM) for filter pruning. Finally, differently from all other APs, BCS-P [56] optimizes the network with a negative log-posterior loss combining a sparsity prior with a robust training objective. Then, the network parameters are sampled from the posterior distribution. For regrowing in D-methods instead, the usual approach is to regrow based on the inverse of the pruning criteria, such as for InTrain [18]. Similarly, TwinRep [55] reparameterizes directly with the product of the two matrices, while BCS-P [56] simply resamples the parameters. FlyingBird [53] instead, uses the Highest Gradient Magnitude (HGM), while DNR [58], following [59], uses normalized momentum.
3.4 Standard vs Adversarial Pruning Methods
After presenting the taxonomy of AP methods, it is fundamental to clarify the difference between standard and AP methods. Following the general pruning formulation in Eq. 1, we identify an objective (i.e., minimizing the loss), and a constraint imposed by pruning to obtain sparsity. While the constraint is the same, the objective clearly changes from standard to AP methods, setting a different goal. Hence, while in standard pruning we will simply minimize the loss $\mathcal{L}$ computed on clean examples, APs consider a robust loss, following Eq. 2:
$\bm{m}^{*}\in\underset{\|\bm{m}\|_{0}\leq k}{\arg\min}\sum_{i=1}^{n}\max_{\|\bm{\delta}_{i}\|_{p}\leq\epsilon}\mathcal{L}(\bm{\theta}\odot\bm{m},\bm{x}_{i}+\bm{\delta}_{i},y)\,,$ (3)
which indicates that the mask $\bm{m}^{*}$ will result from minimizing the loss computed on the adversarial example $\bm{x}+\bm{\delta}$, following Eq. 2. This difference can be encountered, through our taxonomy, by observing the objectives employed in the pipelines, which mostly resort to AT techniques. Similarly, the AP criteria presented in Sect. 3.3.3 explicitly incorporate robustness into their design: e.g., LIS prunes based on the importance of the parameter with respect to adversarial robustness; SOLWM solves the robust constrained optimization problem; RELWM encourages sparsity while enhancing robustness. However, while the objective is discriminant, both standard and AP methods prune weights. Hence, while some details of the pipelines and specifics are unique to the kind of pruning, some other simply abstract from it; e.g., both standard and AP define locality and structure, and considering the sub-optimality of naïve pruning criteria for both tasks, both standard and AP methods can be found to use LWM.
The challenge of AP methods. Finally, let us remark a major challenge of AP methods compared to standard pruning. Considering adversarial training from Eq. 2, and as also highlighted in [12], building a robust model requires designing a much more complex decision boundary. As we prune a model, and reduce the number of parameters, retaining a complex boundary becomes harder. Thus, while performing AT is challenging per se, on a pruned model the challenge escalates, as represented in Figure 2.
Figure 2: A representation of the boundary complexity for different model-task pairs, where red crosses indicate a misclassified or possibly adversarial example. When the model is pruned and the task is standard accuracy, creating a complex decision boundary becomes harder compared to a dense model (first and second figures). When the task is robustness, such difficulty escalates as the boundary is “trained” including adversarial examples within a given perturbation norm (e.g., an $\ell_{\infty}$ constraint).
@@ 4 Benchmarking Adversarial Pruning
In Sect. 3, we classified the AP methods based on pipeline and specifics, and then analyzed the core disparities. Upon a categorization of the methods, we took on the challenge of comparing each AP’s robustness while acknowledging the proposed taxonomy. However, our investigation revealed the impracticability of directly deriving such a comparison due to the many differences in the experimental setups of these evaluations. In fact, newly published APs are compared on equal setups only to a few of the “top-notch” AP methods, and while such analysis can be sufficient to assess the AP novelty, it is not enough to compare the robustness of each AP method (i.e., our goal). Furthermore, we find multiple issues associated with the adversarial evaluations conducted in these works, which represent a well-known problem in the field of adversarial ML, ultimately leading to overestimating the robustness of the models [25, 62]. All these obstacles are implicit limitations in the absence of a benchmark. In this section, we thus present our benchmark implementation and re-evaluate the pruned models produced by AP methods with a comparable experimental setup and fair adversarial setting. In this regard, we first bring to light, in Sect. 4.1, a number of problems related to the current AP evaluations, which are often found to be below par with respect to previous work. We then present the benchmark experimental settings in Sect. 4.2 and re-evaluate the available AP methods in Sect. 4.4, where we conclude by discussing the re-evaluation results with particular attention to our taxonomy.